A experiment calls for 40 gallons of a 70% saline solution. You only have a 80% saline solution and a 40% saline solution. Let x represent the amount of 80% saline solution and y represent the amount of 40% saline solution, what is the equation that describes the total amount of pure saline in the solution?
Answer by addingup(3677)

0.8x + 0.4(40-x) = 0.7(40)
0.8x + 16 - 0.4x = 28
0.4x = 12
x = 30 You need 30 gallons of the 80% solution and 40-30 = 10 gallons of the 40% solution